# Constructing an Optimal Behavior Basis for the Option Keyboard

### v1

@@ 5 Experiments
In this section, we empirically evaluate OKB and investigate the following research questions: Q1: Can OKB approximate a CCS more effectively while requiring fewer base policies than competing methods? Q2: Does OKB’s performance advantage become more pronounced as problem complexity increases, i.e., as the number of reward features $d$ grows? Q3: Can the base policies learned by OKB be leveraged to solve tasks with non-linear reward functions under the conditions outlined in Thm. 4.5?
@@
Fig. 1 depicts the domains used in our experiments. To handle the high-dimensional state space of these domains, we learn base policies using a Universal SF Approximator (USFA) Borsa et al., (2019). Furthermore, rather than learning a separate SF for each base policy $\pi\in\Pi_{k}$, we train a single USFA, $\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})$, conditioned on task vectors $\bm{\mathbf{w}}$, such that $\pi_{\bm{\mathbf{w}}}(s)\approx\operatorname*{arg\,max}_{a\in\mathcal{A}}\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})\cdot\bm{\mathbf{w}}$. At each call to NewPolicy($\bm{\mathbf{w}}$) in Alg.1, the USFA is trained to solve the new task $\bm{\mathbf{w}}$. Additional experimental details are provided in App. E.
(a) Minecart domain.
(b) Highway domain.
Figure 2: Mean normalized return per iteration for each method on a set of test tasks in the (a) Minecart and (b) Highway domains.
Figure 3: Mean normalized return over test tasks as a function of iteration number—i.e., number of base policies learned per method [FetchPickAndPlace]. As $d$ increases, OKB’s performance advantage over SFOLS (a state-of-the-art GPI-based algorithm) grows.
@@
In all experiments, we report the mean normalized return of each method (normalized with respect to the minimum and maximum returns observed for a given task) along with the 95% bootstrapped confidence interval over 15 random seeds. We compare OKB to SFOLS (Alegre et al., , 2022), a method that identifies a CCS using GPI policies, and to OKB-Uniform, a variant of OKB that selects tasks uniformly from $\mathcal{W}$ in line 21 of Alg. 1, rather than from $\mathcal{C}$. At each iteration, all methods have a fixed budget of environment interactions. To ensure fair comparisons, since OKB must also train a meta-policy while SFOLS does not, we restrict OKB to using only half of its budget for learning base policies, allocating the other half to training the meta-policy (Alg. 2). This makes the comparison more conservative for OKB, as it has fewer interactions available to learn base policies.
@@
In Fig.2, we show the mean return of each method over a set of test tasks, $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, [^4] To generate test task sets $\mathcal{W}^{\prime}\subset\mathcal{W}$ for different values of $d$, we used the method introduced by Takagi et al., (2020), which produces uniformly spaced weight vectors in $\mathcal{W}$. as a function of the iteration number (i.e., the number of base policies learned). We report results for the Minecart domain—a classic multi-objective RL problem—and the Highway domain. Both domains have reward functions defined by $d=3$ reward features, which are detailed in App. E.2. These results demonstrate that OKB achieves strong performance across test tasks with a small number of base policies, positively answering research question Q1: OKB can approximate a CCS more effectively using fewer base policies. This is particularly evident in Fig. 2(a), where OKB reaches near-optimal performance with just 2–3 base policies. While OKB-Uniform also performs well in the Minecart domain, its lower performance in the Highway domain (Fig. 2(b)) highlights the importance of expanding the set of base policies by carefully selecting promising tasks—defined by corner weights—for training. Across both domains, OKB consistently outperforms all competing methods.
@@
To investigate Q2, we evaluate each method in the FetchPickAndPlace domain with varying numbers of reward features, $d\in\{2,4,6,8\}$. In this domain, an agent controls a robotic arm that must grasp a block from a table and move it to a specified location. Each of the $d$ reward features represents the block’s distance to a different target location (shown in red in Fig. 1). Fig. 3 shows that as the number of target locations ($d$) increases, the performance gap between OKB and SFOLS increases significantly—positively answering Q2. Intuitively, OKB focuses on tasks where the OK cannot express optimal actions (see Thm. 4.2). By learning the corresponding optimal policies, OKB quickly identifies a behavior basis that enables the OK to solve tasks across the entire space of task vectors $\mathcal{W}$. Conversely, the gap between OKB and OKB-Uniform decreases since a larger number of reward features enhances the expressivity of the OK (Prop. 4.1). The total computation time required by OKB to train a meta-policy and base policies to ensure zero-shot optimality is consistently comparable to or lower than that of competing methods. This is because OKB requires fewer policies to be learned and evaluated at inference time.
Figure 4: Mean total reward in the Item Collection domain under a non-linear reward function.
Figure 5: Continuous actions produced by the meta-policy while solving a sample task. The color of each timestep’s column indicates the selected base policy. Notice OKB selects base policies in a temporally consistent manner, suggesting these policies encode temporally-extended behaviors useful to solving the task.
@@
Next, we investigate Q3 by conducting an experiment similar to the one proposed by Alver and Precup, (2022). We compare OKB to relevant competitors in an environment with non-linear reward functions. The Item Collection domain (Fig. 1; top right) is a $10\times 10$ grid world where agents must collect two types of items. Reward features are indicator functions that signal whether the agent has collected a particular item in the current state. This domain requires function approximation due to the combinatorial number of possible states. After OKB learns a behavior basis $\Pi_{k}$, the agent trains a meta-policy $\omega:\mathcal{S}\to\mathcal{Z}$ to solve a task with a non-linear reward function. Specifically, this is a sequential task where the agent must collect all instances of one item type before collecting any items of another type. We compare OKB with SIP (Alver and Precup, , 2022), [^5] SIP was not included in previous experiments since it assumes independent reward features (Alver and Precup, , 2022). a method that learns $d$ base policies—each maximizing a specific reward feature—and a meta-policy $\omega$ over a discrete set of weights in $\mathcal{W}$. We also compare OKB against a baseline DQN agent that learns tasks from scratch and against GPI policies (horizontal blue lines), which are optimal for either exclusively prioritizing one item type or assigning equal importance to both. Fig. 4 shows the mean total reward achieved by each method in this non-linear task as a function of the total number of environment interactions required to train the meta-policies. Both OKB and SIP successfully solved the task, with OKB doing so more quickly. The DQN baseline failed to solve the task as directly learning a policy over $\mathcal{A}$ is significantly more difficult.
@@
Finally, we examine the qualitative behavior of OKB policies by visualizing the continuous actions produced by the meta-policy $\omega(s,\bm{\mathbf{w}})$ while solving a sample task in the Highway domain, where the agent must prioritize driving fast and staying in the rightmost lane. In Fig. 5, the color of each timestep’s column represents the selected base policy. The agent initially accelerates as quickly as possible by following base policy $\pi_{1}$ for the first 10 timesteps while the rightmost lane is occupied. At timestep $t=10$, the agent turns right and switches to base policy $\pi_{2}$, which is optimized for driving in the rightmost lane. At $t=20$, $\pi_{1}$ becomes active again, allowing the agent to accelerate while staying in the rightmost lane. Notice that while the meta-policy $\omega$ is relatively smooth, the base policies it selects are complex. In particular, although $\bm{\mathbf{z}}_{t}$ remains approximately constant during the first 10 timesteps, the underlying primitive actions $a\in\mathcal{A}$ (Fig. 5; black numbers at the top of each column) continuously alternate between turning left, turning right, accelerating, and idling. Finally, notice that OKB selects base policies in a temporally consistent manner, suggesting that it learns to identify temporally extended behaviors—akin to options Sutton et al., (1999)—that help solve the task.
@@ 6 Related Work
OK and GPI.
Previous works have extended the OK in different ways. Carvalho et al., 2023b introduced a neural network architecture for jointly learning reward features and SFs, while Machado et al., (2023) proposed using the OK with base policies that maximize Laplacian-based reward features. However, unlike our work, these methods do not provide theoretical guarantees on optimality for solving specific families of tasks. Other works have introduced GPI variants for risk-aware transfer (Gimelfarb et al., , 2021), mitigating function approximation errors (Kim et al., , 2024), planning with approximate models (Alegre et al., 2023a, ), and combining non-Markovian policies (Thakoor et al., , 2022). These approaches improve GPI in ways that are orthogonal to our contribution and could potentially be combined with our method.
Learning behavior basis.
Previous works have addressed the problem of learning an effective behavior basis for transfer learning within the SF framework (Nemecek and Parr, , 2021; Zahavy et al., , 2021). The method introduced by Alver and Precup, (2022) assumes that reward features are independent (i.e., can be controlled independently) and that the MDP’s transition function is deterministic. These assumptions often do not hold in complex RL domains, including most of those studied in Section 5. Alegre et al., (2022) proposed a method that learns a set of policies corresponding to a CCS and combines them with GPI. In contrast, we use option keyboards, which enable a broader range of policies to be expressed, allowing our method to approximate a CCS more effectively with a smaller behavior basis.
Learning features.
While we focused on identifying an optimal behavior basis with respect to a given set of reward features, other works have addressed the complementary problem of learning more expressive reward features (Touati and Ollivier, , 2021; Carvalho et al., 2023a, ; Chua et al., , 2024). A promising future research direction is to combine OKB with methods such as the Forward-Backward representation Touati et al., (2023) to construct an optimal behavior basis under learned reward features.
@@ 7 Conclusions
We introduced OKB, a principled method with strong theoretical guarantees for identifying the optimal behavior basis for the Option Keyboard (OK). Our theoretical results, supported by thorough empirical analysis, show that OKB significantly reduces the number of base policies required to achieve zero-shot optimality in new tasks compared to state-of-the-art methods. In particular, OKB constructs a behavior basis efficiently by carefully selecting base policies that iteratively improve its meta-policy’s approximation of the CCS. We empirically evaluate OKB in challenging high-dimensional RL problems, demonstrating that it consistently outperforms GPI-based approaches. Notably, its performance advantage becomes increasingly pronounced as the number of reward features grows. Finally, we prove that OKB’s expressivity surpasses that of a CCS, enabling it to optimally solve specific classes of non-linear tasks.
An interesting direction for future research is extending OKB with temporally extended meta-policies $\omega$ that incorporate learned termination conditions. Additionally, given the recently identified mathematical connections between multi-objective RL and SFs (Alegre et al., , 2022), OKB could be adapted and combined with existing MORL techniques to further improve sample complexity.
@@ Appendix A Proofs
Let $\Pi_{k}$ be a set of base policies and $\omega^{*}$ be a meta-policy trained to convergence to solve a given task $r$. If $A^{\omega^{*}}{r}(s,a)>0$ for some $(s,a)$, then action $a$ is not expressible by $\pi^{\text{OK}}{\omega^{*}}(\cdot;\Pi_{k})$. Consequently, the OK with base policies $\Pi_{k}$ cannot represent an optimal policy for $r$.
@@
If $A^{\omega}_{r}(s,a)>0$, it means executing $a$ in $s$ provides a higher expected return than the action currently selected by $\pi^{\text{OK}}_{\omega}$. Hence, if there exists an $(s,a)$ such that $A^{\omega}_{r}(s,a)>0$, it implies that the OK policy fails to represent at least one optimal action, meaning it is suboptimal.
@@
Since $\omega^{*}$ is trained until convergence to solve $r$, it has reached a stable policy. However, if there exists $(s,a)$ such that $A^{\omega^{*}}_{r}(s,a)>0$, then:
@@
This means that in at least one state-action pair, the OK policy does not select an action that maximizes the expected return.
Since $\pi^{\text{OK}}_{\omega^{*}}$ is restricted to the actions expressible via its base policies $\Pi_{k}$, this implies that:
@@
Since at least one optimal action $a^{*}$ is missing from the action space of the OK policy, it follows that: $\pi^{OK}_{\omega^{*}}$ is not an optimal policy for $r$, and the task $r$ cannot be optimally solved using the given base policies $\Pi_{k}$. Therefore, to guarantee optimality, additional base policies must be introduced into $\Pi_{k}$. ∎
@@
The result in this lemma follows directly from the theoretical guarantees of previous methods that construct a CCS, e.g., OLS and SFOLS (Roijers, , 2016; Alegre et al., , 2022). Intuitively, this result states that to check if a given SF vector set $\Psi_{k}$ is a CCS, it is enough to check if it contains an optimal policy for all of its corner weights. If this is the case, then it is not possible to identify some alternative weight vector $\bm{\mathbf{w}}\in\mathcal{W}$ for which it does not contain an optimal policy.
@@
OKB returns a set of base policies $\Pi_{k}$ and a meta-policy $\omega$ such that $|\Pi_{k}|\leq|\mathrm{CCS}|$ and $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is optimal for any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. This implies that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$; i.e., the OK is at least as expressive as a CCS.
@@
Assumption 1. NewPolicy($\bm{\mathbf{w}}$) returns an optimal policy $\pi_{\bm{\mathbf{w}}}^{*}$ for the given task $\bm{\mathbf{w}}$.
Assumption 2. For all $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$, if an optimal policy $\pi_{\bm{\mathbf{w}}}^{*}\in\Pi^{\text{OK}}(\Pi)$, then TrainOK($\omega,\mathcal{W}^{\text{sup}},\Pi$) (Alg. 3) returns a meta-policy $\omega$ s.t. $\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi)$ is optimal w.r.t. task $\bm{\mathbf{w}}$.
OKB (Alg. 1) maintains two partial CCSs, $\Psi^{\text{base}}_{k}$ and $\Psi^{\text{OK}}_{k}$, respectively being the SFs of the base policies $\Pi_{k}$ and the policies generated via the OK using $\Pi_{k}$. We first will prove that, if in any given iteration $k$, the set of candidate corner weights $\mathcal{C}$ is empty (see line 16 in Alg. 1), then it holds that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}_{k+1}$. First, note that $\mathcal{C}$ contains corner weights of both $\Psi^{\text{base}}_{k}$ and $\Psi^{\text{OK}}_{k}$. If the OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is able to optimally solve all tasks in $\mathcal{C}$ (line 12 in Alg. 1), then we have that $\mathrm{CCS}\in\Psi^{\text{OK}}(\Pi_{k})$ due to Lemma A.1. We note that assumptions 1 and 2 above are necessary because Lemma A.1 requires the partial CCS $\Psi_{k}$ to be a subset of the CCS. If it contains $\epsilon$-optimal policies, then it is possible to prove convergence to $\epsilon$-CCSs instead (Alegre et al., 2023b, ).
Now, to conclude the proof, we only need to show that $|\Pi_{k}|\leq|\mathrm{CCS}|$. The set of base policies is never larger than the CCS due to line 24 in Algorithm 1, in which dominated base policies (i.e., redundant policies that are not exclusively optimal w.r.t. any $\bm{\mathbf{w}}$) are removed from $\Pi_{k}$ by the procedure RemoveDominated. Importantly, the set of base policies returned by OKB, $\Pi_{k}$, will only have the same of the CCS ($|\Pi_{k}|=|\mathrm{CCS}|$) in domains where no single task in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ can be optimally solved by combining policies optimal for other tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$.  [^6] An example of such a domain is an MDP composed of independent corridors, in which optimal policies for different weights $\bm{\mathbf{w}}\in\mathcal{W}$ must traverse different corridors.
A.4 Proof of Theorem. 4.5
Theorem (4.5).

### v2

@@ 5 Experiments
We now empirically evaluate OKB and investigate the following research questions: Q1: Can OKB approximate a CCS more effectively while requiring fewer base policies than competing methods? Q2: Does OKB’s performance advantage grow with problem complexity, as measured by the number of reward features $d$? Q3: Can the base policies learned by OKB be used to solve tasks with non-linear reward functions, under the conditions in Prop. 4.4?
@@
Fig. 1 depicts the domains used in our experiments. To handle the high-dimensional state space of these domains, we learn base policies using a Universal SF Approximator (USFA) (Borsa+2019). Furthermore, rather than learning a separate SF for each base policy $\pi\in\Pi_{k}$, we train a single USFA, $\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})$, conditioned on task vectors $\bm{\mathbf{w}}$, such that $\pi_{\bm{\mathbf{w}}}(s)\approx\operatorname*{arg\,max}_{a\in\mathcal{A}}\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})\cdot\bm{\mathbf{w}}$. Each call to NewPolicy($\bm{\mathbf{w}}$) in Alg.1 trains the USFA to solve task $\bm{\mathbf{w}}$. We employ USFAs in our experiments primarily to avoid the memory and computational costs of training a separate neural network for each policy’s SFs. Moreover, USFAs typically achieve a level of approximation comparable to—or better than—that of independently trained networks, since in both cases generalization depends mainly on the architecture, capacity, and distribution of training data. See App. F for more details.
@@
In all experiments, we report the mean normalized return of each method (normalized with respect to the minimum and maximum returns observed for a given task) along with the 95% bootstrapped confidence interval over 15 random seeds. We compare OKB to SFOLS (Alegre+2022), a method that identifies a CCS using GPI policies, and to OKB-Uniform, a variant of OKB that selects tasks uniformly from $\mathcal{W}$ in line 14 of Alg. 1, rather than from $\mathcal{C}$. An iteration corresponds to each method training a base policy within a fixed budget of environment interactions. To ensure fair comparisons, since OKB must also train a meta-policy while SFOLS does not, we restrict OKB to using only half of its budget for learning base policies, allocating the other half to training the meta-policy (Alg. 2). This makes the comparison more conservative for OKB, as it has fewer interactions available to learn base policies.
(a) Minecart
(b) Highway
Figure 2: Mean normalized return per iteration for each method on a set of test tasks.
@@
Fig. 2 depicts each method’s mean return over a set of test tasks, $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, [^4] To generate test task sets $\mathcal{W}^{\prime}\subset\mathcal{W}$ for different values of $d$, we used the method introduced by Takagi+2020, which produces uniformly spaced weight vectors in $\mathcal{W}$. as a function of the iteration number (i.e., the number of base policies learned). We report results for the Minecart domain—a classic multi-objective RL problem—and the Highway domain. Both domains have reward functions defined by $d=3$ reward features, which are detailed in App. F.2. These results demonstrate that OKB achieves strong performance across test tasks with a small number of base policies, positively answering research question Q1: OKB can approximate a CCS more effectively using fewer base policies. This is particularly evident in Fig. 2(a), where OKB reaches near-optimal performance with just 2–3 base policies. While OKB-Uniform also performs well in the Minecart domain, its lower performance in the Highway domain (Fig. 2(b)) highlights the importance of expanding the set of base policies by carefully selecting promising tasks—defined by corner weights—for training. Across both domains, OKB consistently outperforms all competing methods.
@@
To investigate Q2, we evaluate each method in the FetchPickAndPlace domain with varying numbers of reward features, $d\in\{2,4,6,8\}$. In this domain, an agent controls a robotic arm that must grasp a block and move it to a specified location. Each of the $d$ reward features represents the block’s distance to a different target location (red spheres in Fig. 1). Fig. 3 shows that as the number of target locations ($d$) increases, the performance gap between OKB and SFOLS increases significantly—positively answering Q2.  [^5] Importantly, to the best of our knowledge, most state-of-the-art methods are typically evaluated on problems with dimensionality at most $d=4$ (e.g., SFOLS and SIP). In our experiments, by contrast, we double the number of reward features, thus showing that our method remains effective and continues to outperform all baselines even on problems an order of magnitude more challenging. Intuitively, OKB focuses on tasks where the OK cannot express optimal actions (see Thm. 4.2). By learning the corresponding optimal policies, OKB quickly identifies a behavior basis that enables the OK to solve tasks across the entire space of task vectors $\mathcal{W}$. Conversely, the gap between OKB and OKB-Uniform decreases since a larger number of reward features enhances the expressivity of the OK (Prop. 4.1). The total computation time required by OKB to train a meta-policy and base policies for zero-shot optimality is consistently comparable to or lower than that of competing methods, as it learns and evaluates fewer policies at inference time.
@@
Next, we investigate Q3 by conducting an experiment similar to the one proposed by Alver&Precup2021. We compare OKB to relevant competitors in an environment with non-linear reward functions. The Item Collection domain (Fig. 1; top right) is a $10\times 10$ grid world where agents must collect two types of items. Reward features are indicator functions that signal whether the agent has collected a particular item in the current state. This domain requires function approximation due to the combinatorial number of possible states. After OKB learns a behavior basis $\Pi_{k}$, the agent trains a meta-policy $\omega:\mathcal{S}\to\mathcal{Z}$ to solve a task with a non-linear reward function. Specifically, this is a sequential task where the agent must collect all instances of one item type before collecting any items of another type. We compare OKB with SIP (Alver&Precup2021), [^6] SIP was not part of previous experiments as it assumes independent reward features (Alver&Precup2021). a method that learns $d$ base policies—each maximizing a specific reward feature—and a meta-policy $\omega$ over a discrete set of weights in $\mathcal{W}$. We also compare OKB against a baseline DQN agent that learns tasks from scratch and against GPI policies (horizontal blue lines), which are optimal for either exclusively prioritizing one item type or assigning equal importance to both. Fig. 5 shows the mean total reward achieved by each method in this non-linear task as a function of the total number of environment interactions required to train the meta-policies. Both OKB and SIP successfully solved the task, with OKB doing so more quickly. The DQN baseline failed to solve the task as directly learning a policy over $\mathcal{A}$ is significantly more difficult.
Figure 3: Mean normalized return over test tasks as a function of iteration number (i.e., number of base policies learned per method) [FetchPickAndPlace]. As $d$ increases, OKB’s performance advantage over SFOLS (a state-of-the-art GPI-based algorithm) grows.
Figure 4: Mean return in the Item Collection domain under a non-linear reward function.
Figure 5: Continuous actions from the meta-policy on a sample task. OKB selects base policies in a temporally consistent way.
@@
Finally, we examine the qualitative behavior of OKB policies by visualizing the continuous actions produced by the meta-policy $\omega(s,\bm{\mathbf{w}})$ while solving a sample task in the Highway domain, where the agent must prioritize driving fast and staying in the rightmost lane. In Fig. 5, the color of each timestep’s column represents the selected base policy. The agent initially accelerates as quickly as possible by following base policy $\pi_{1}$ for the first 10 timesteps while the rightmost lane is occupied. At timestep $t=10$, the agent turns right and switches to base policy $\pi_{2}$, which is optimized for driving in the rightmost lane. At $t=20$, $\pi_{1}$ becomes active again, allowing the agent to accelerate while staying in the rightmost lane. Note that while the meta-policy $\omega$ is relatively smooth, the base policies it selects are complex. In particular, although $\bm{\mathbf{z}}_{t}$ remains approximately constant during the first 10 timesteps, the underlying primitive actions $a\in\mathcal{A}$ (Fig. 5; black numbers at the top of each column) continuously alternate between turning left or right, accelerating, and idling. Finally, notice that OKB selects base policies in a temporally consistent manner, suggesting that it learns to identify temporally extended behaviors—akin to options (Sutton+1999)—that help solve the task.
@@ 6 Related Work
OK and GPI. Previous works have extended the OK in different ways. Carvalho+2023neurips introduced a neural network architecture for jointly learning reward features and SFs, while Machado+2023 proposed using the OK with base policies that maximize Laplacian-based reward features. However, unlike our work, these methods do not provide theoretical guarantees on optimality for solving specific families of tasks. Other works have introduced GPI variants for risk-aware transfer (Gimelfarb+2021), mitigating function approximation errors (Kim+2022neurips), planning with approximate models (Alegre+2023neurips), planning to solve tasks defined by finite state automata (Kuric+2024), and combining non-Markovian policies (Thakoor+2022). These approaches improve GPI in ways that are orthogonal to our contribution and could potentially be combined with our method.
Learning behavior basis. Previous works have addressed the problem of learning an effective behavior basis for transfer learning within the SF framework (Nemecek&Parr2021; Zahavy+2021). The method introduced by Alver&Precup2021 assumes that reward features are independent (i.e., can be controlled independently) and that the MDP’s transition function is deterministic. These assumptions often do not hold in complex RL domains, including most of those studied in Section 5. Alegre+2022 proposed a method that learns a set of policies corresponding to a CCS and combines them with GPI. In contrast, we use option keyboards, which enable a broader range of policies to be expressed, allowing our method to approximate a CCS more effectively with a smaller behavior basis.
Learning features. While we focused on identifying an optimal behavior basis given a particular set of reward features, other works have addressed the complementary problem of learning more expressive reward features (Touati&Ollivier2021; Carvalho+2023; Chua+2024). A promising future research direction is to combine OKB with methods such as the Forward-Backward representation (Touati+2023) to construct an optimal behavior basis under learned reward features.
@@ 7 Conclusions
We introduced OKB, a principled method with strong theoretical guarantees for identifying the optimal behavior basis for the Option Keyboard (OK). Our theoretical results, supported by thorough empirical analysis, show that OKB significantly reduces the number of base policies required to achieve zero-shot optimality in new tasks compared to state-of-the-art methods. In particular, OKB constructs a behavior basis efficiently by carefully selecting base policies that iteratively improve its meta-policy’s approximation of the CCS. We empirically evaluate OKB in challenging high-dimensional RL problems, demonstrating that it consistently outperforms GPI-based approaches. Notably, its performance advantage becomes increasingly pronounced as the number of reward features grows. Finally, we prove that OKB’s expressivity surpasses that of a CCS, enabling it to optimally solve specific classes of non-linear tasks. Among several possible research directions, one promising direction is to extend OKB with temporally extended meta-policies that incorporate learned termination conditions. The resulting expressiveness could further expand OKB’s flexibility and efficiency in reusing behaviors across tasks. Furthermore, recently identified mathematical connections between multi-objective RL (MORL) and SFs (Alegre+2022) open the door to adapting OKB to integrate with MORL techniques, paving the way for unified representations that improve both sample efficiency and generalization across objectives.
Acknowledgments
This work is partially supported by CAPES (Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brazil, Finance Code 001). Ana Bazzan is partially supported by CNPq under grant number 304932/2021-3.
Appendix
@@ Appendix A Proofs
Let $\Pi_{k}$ be a set of base policies and $\omega^{*}$ be a meta-policy trained to convergence to solve a given task $r$. If $A^{\omega^{*}}_{r}(s,a)>0$ for some $(s,a)$, then action $a$ is not expressible by $\pi^{\text{OK}}_{\omega^{*}}(\cdot;\Pi_{k})$. Consequently, the OK with base policies $\Pi_{k}$ cannot represent an optimal policy for $r$.
@@
If $A^{\omega}_{r}(s,a)>0$, then executing $a$ in $s$ provides a higher expected return than the action currently selected by $\pi^{\text{OK}}_{\omega}$. Hence, if there exists an $(s,a)$ such that $A^{\omega}_{r}(s,a)>0$, this implies that the OK policy fails to represent at least one optimal action, meaning it is suboptimal.
@@
Since $\omega^{*}$ is trained until convergence to solve $r$, it has reached a stable policy. However, if there exists an $(s,a)$ such that $A^{\omega^{*}}_{r}(s,a)>0$, then
@@
This means that in at least one state-action pair, the OK policy does not select an action that maximizes the expected return. Since $\pi^{\text{OK}}_{\omega^{*}}$ is restricted to the actions expressible via its base policies $\Pi_{k}$, this implies that
@@
Since at least one optimal action $a^{*}$ is missing from the action space of the OK policy, it follows that $\pi^{OK}_{\omega^{*}}$ is not an optimal policy for $r$, and the task $r$ cannot be optimally solved using the given base policies $\Pi_{k}$. Therefore, to guarantee optimality, additional base policies must be introduced into $\Pi_{k}$. ∎
@@
This lemma follows directly from the theoretical guarantees of previous methods that construct a CCS, e.g., OLS and SFOLS (Roijers2016; Alegre+2022). Intuitively, this result states that to check if a given set of SF vectors, $\Psi_{k}$, is a CCS, it is sufficient to check if it contains an optimal policy for all of its corner weights. If this is the case, then it is not possible to identify some alternative weight vector $\bm{\mathbf{w}}\in\mathcal{W}$ for which it does not contain an optimal policy.
@@
OKB returns a set of base policies, $\Pi_{k}$, and a meta-policy, $\omega$, such that $|\Pi_{k}|\leq|\mathrm{CCS}|$ and $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is optimal for any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. This implies that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$; i.e., the OK matches or exceeds the expressiveness of a CCS.
@@
We start the proof by stating two assumptions on which the theorem relies:
Assumption 1. NewPolicy($\bm{\mathbf{w}}$) returns an optimal policy, $\pi_{\bm{\mathbf{w}}}^{*}$, for the given task $\bm{\mathbf{w}}$.
Assumption 2. For all $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$, if an optimal policy $\pi_{\bm{\mathbf{w}}}^{*}\in\Pi^{\text{OK}}(\Pi)$, then TrainOK($\omega,\mathcal{W}^{\text{sup}},\Pi$) (Alg. 3) returns a meta-policy, $\omega$, s.t. $\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi)$ is optimal w.r.t. task $\bm{\mathbf{w}}$.
Recall that the OKB (Alg. 1) maintains two partial CCSs, $\Psi^{\text{base}}_{k}$ and $\Psi^{\text{OK}}_{k}$, which store, respectively, the SFs of the base policies $\Pi_{k}$ and the policies generated via the OK using $\Pi_{k}$. We first will prove that, if in any given iteration $k$, the set of candidate corner weights $\mathcal{C}$ is empty (see line 16 in Alg. 1), then it holds that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}_{k+1}$. First, note that $\mathcal{C}$ contains corner weights of both $\Psi^{\text{base}}_{k}$ and $\Psi^{\text{OK}}_{k}$. If the OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is able to optimally solve all tasks in $\mathcal{C}$ (line 12 in Alg. 1), then we have that $\mathrm{CCS}\in\Psi^{\text{OK}}(\Pi_{k})$ due to Lemma A.1. Note that assumptions 1 and 2, above, are necessary because Lemma A.1 requires the partial CCS $\Psi_{k}$ to be a subset of the CCS. If it contains $\epsilon$-optimal policies, then it is possible to prove convergence to $\epsilon$-CCSs instead (Alegre+2023).
Now, to conclude the proof, we only need to show that $|\Pi_{k}|\leq|\mathrm{CCS}|$. The set of base policies is never larger than the CCS due to line 24 in Alg. 1, in which dominated base policies (i.e., redundant policies that are not exclusively optimal w.r.t. any $\bm{\mathbf{w}}$) are removed from $\Pi_{k}$ by the procedure RemoveDominated. Importantly, the set of base policies returned by OKB, $\Pi_{k}$, will only have the same of the CCS ($|\Pi_{k}|=|\mathrm{CCS}|$) in domains where no single task in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ can be optimally solved by combining policies optimal for other tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. [^7] An example of such a domain is an MDP composed of independent corridors, in which optimal policies for different weights $\bm{\mathbf{w}}\in\mathcal{W}$ must traverse different corridors.
A.4 Proof of Proposition 4.4
Proposition (4.4).